Hi there hows it going? I was wondering if you would be able to assist me please. I am stuck on how to work out a word problem, regarding a rectangle. It goes like this: The perimeter of a rectangular field is 184m, and its area is 1920m^2. Find the dimensions of the recangle.
I have worked out that the dimensions are in fact 32m x 60m = 1920m^2, or (32 x 2) + (60 x 2) = 184m. However what I need to know is how to work this out with a formular as I will not always have an entire hour to muck around. Your help would be greatly appreciated. Cheers D.

Answer by edjones(8007)

2L+2w=184
L*w=1920
.
L=1920/w
2(1920/w)+2w=184 Substitute 1920/w for L
3840/w+2w=184
3840+2w^2=184w multiply each side by x to eliminate fractions.
2w^2-184w+3840=0
2(w^2-92+1920)=0
w^2-92+2116=-1920+2116 complete the square
(w-46)^2=196
w-46=+-14
w=46+-14
w=60, w=32
.
Ed